evaluate the following integral
∫ sec x cosec x / log(tan x) dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello student,
Please find the answer to your question below

I = \int \frac{secxcosecx}{log(tanx)}dx
log(tanx) = t
\frac{1}{tanx}sec^{2}xdx = dt
secxcosecxdx = dt
I = \int \frac{1}{t}dt
I = log(t)+c
I = log(log(tanx))+c